The points (0, 8/3),
(1,3) and (82, 30) are vertices of


(a) isoceles triangle


(b) an acute angled triangle


(c) right angled triangle


(d) None of these


Explain pls.

Grade:12th Pass

3 Answers

Aman Bansal
12 years ago

Dear Aditi,

square of one side ,a^2 = { (0-1)^2 + (8/3 -3)^2  } = 10/9 =1.11

b^2 = { (82-1)^2 + (30 -3)^2  } = 7290

 c^2 =  { (82-0)^2 + (30 -8/3)^2  } =7471.11

it shows that , c^2 > a^2 + b^2

so the tiangle would be acute

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

Aditya naitan
7 years ago
If we will find the area of triangle it would be 0. therefore the points would be collinear. the correct option is “none of these”.
Abcd
6 years ago
Find area of triangle which comes out to be zero. Hence the points are collinear. So option D is correct